OpenStudy (anonymous):

36. How many milliliters of juice are in a 2-liter bottle of juice? A. 2000 mL B. 1000 mL C. 200 mL D. 20 mL

OpenStudy (anonymous):

A

OpenStudy (goldphenoix):

A. 2000 ML. 1 L = 1000 ML. I hope this help!

OpenStudy (dan815):

HOPE SO

OpenStudy (anonymous):

Bam I said it first GOLD.

OpenStudy (goldphenoix):

You didn't prove your answer. :) I proved my answer. :P

OpenStudy (anonymous):

Whatever, I still got it right.

OpenStudy (anonymous):

ickbeth, i dont get a medal? ;(

OpenStudy (anonymous):

mikeyy1992 you might have said it first but you didn't explain it

hero (hero):

^What @ickbeth said

OpenStudy (anonymous):

.-. k.

OpenStudy (jhannybean):

lol.

OpenStudy (anonymous):

A. 2000 mL So you are left with this for the answer. B. 1000 mL This amount is equal to one liter, but the question asks for two. C. 200 mL This is only 1/5 of a liter. D. 20 mL It wouldn't be this, because this is too little to even be one liter.

hero (hero):

x = amount to be found \(x \text{mL} = 2L\) m = 1/1000 \(x \frac{1}{1000}L = 2L\) Cancel L on both sides: \(x \frac{1}{1000} = 2\) \(\frac{x}{1000} = 2\) Multiply both sides by 1000 x = 2000

OpenStudy (goldphenoix):

Oh please, stop trying. >_> I already got the medal. :)

OpenStudy (mathstudent55):

milli- = 1/1000 1 milliliter = 1 mL = (1/1000) liter 1000 mL = 1 liter

OpenStudy (anonymous):

I now have a grudge with new people on here.